# HOFT: Householder Orthogonal Fine-tuning

### v1

@@ 4.2 Machine Translation
For measuring machine translation performance, HOFT and SHOFT are compared with DoRA and LoRA using four languages from the CoVoST 2 [47] dataset: Slovene, German, Latvian and French. We chose these languages in order to have two well-represented languages and two low-resource languages. For French and German, models are trained on the first 10K elements of the training split. Three models are adapted for this task: NLLB-3.3B [11], LLaMA2-7B [45], and LLaMA3.1-8B [16]. We set $r=16$ for all PEFT methods and train the models for 2 epochs. Both BLEU [33, 35] and COMET [38, 39] results are provided for each individual language per model and PEFT method. Results obtained are shown in Table 3. We additionally provide baseline performance of the models.
[…]
Model Method #Params (%) Slovene German Latvian French
BLEU COMET BLEU COMET BLEU COMET BLEU COMET
NLLB-3.3B Baseline - 39.7 87.5 39.3 86.2 31.2 81.3 38.5 84.9
LoRA 0.42 46.8 89.2 44.5 87.7 38.2 83.9 49.7 87.8
DoRA 0.43 46.8 89.1 44.7 87.6 38.2 83.9 49.5 87.7
HOFT 0.42 48.0 89.4 44.4 87.6 38.6 83.9 49.5 87.7
SHOFT 0.43 46.4 89.5 44.5 87.7 38.7 84.0 49.7 87.8
LLaMA2-7B 0-shot - 26.8 72.8 30.4 74.1 4.5 52.2 37.2 79.3
LoRA 0.19 39.3 84.7 41.5 86.9 15.5 66.2 47.0 87.2
DoRA 0.19 39.6 84.8 41.4 86.9 16.2 66.6 47.0 87.2
HOFT 0.19 40.6 85.2 41.4 86.9 15.8 66.6 47.0 87.3
SHOFT 0.19 41.2 85.6 41.6 87.0 15.7 65.9 47.1 87.3
LLaMA3.1-8B 0-shot - 34.2 77.8 40.9 86.2 22.9 70.8 41.6 82.7
LoRA 0.12 36.2 84.1 42.3 87.4 32.7 80.9 46.8 85.5
DoRA 0.12 42.4 85.0 42.2 87.4 32.8 80.8 46.7 85.5
HOFT 0.12 44.2 86.6 42.9 87.5 32.2 80.4 46.7 85.6
SHOFT 0.12 43.6 86.4 43.1 87.7 31.9 80.4 46.8 85.6
From Table 3 we can observe how HOFT and SHOFT provide competitive results in French and German. In Latvian, HOFT and SHOFT give similar results in the case of NLLB-3.3B. For Slovene, both methods clearly outperform LoRA and DoRA with LLaMA2-7B, while HOFT in BLEU and SHOFT in COMET with NLLB-3.3B. Notably, the difference on both metrics is significantly higher with LLaMA3.1-8B. Overall, the top BLEU and COMET scores are almost always achieved by either HOFT or SHOFT, underlining their effectiveness across multiple languages.
@@ 4.3 Subject-driven generation
For subject-driven generation, we follow the experimental protocol of HRA [53], using the DreamBooth dataset [41] to train and evaluate on 25 distinct subjects, each with 30 associated prompts. We adapt the pre-trained Stable Diffusion (SD) model [40] and compare PEFT methods quantitatively across four metrics: subject fidelity (DINO [7] and CLIP-I [37]), prompt fidelity (CLIP-T [37]), and sample diversity (LPIPS [56]).
@@
The results, together with the provided baselines, are summarized in Table 4. Both HOFT and SHOFT outperform all baselines in subject fidelity. In terms of textual prompt fidelity, they achieve results comparable with LoRA, OFT, and COFT. For sample diversity, they also deliver competitive performance. Additionally, we also tested HOFT and SHOFT at half the rank. Even with fewer trainable parameters, both methods consistently outperform LoRA, OFT, and COFT across all metrics, while remaining competitive with HRA on subject fidelity.
@@
Therefore, in order to gain a deeper insight into subject fidelity, we conducted an additional experiment following DoRA [27]. We fine-tuned a pre-trained Stable Diffusion XL (SDXL) model [34] on two datasets: 3D icons and lego sets. In Figure 3 we can see some examples of the styles to be learned. In this experiment, five PEFT methods are used for evaluation: LoRA, HRA, OFT, HOFT, and SHOFT. To ensure a fair comparison, all methods used the same random sample seed for generating the images.
As shown in Figure 4, HOFT and SHOFT provide better personalization than LoRA, HRA, and OFT. When generating 3D icons, both methods closely match the style and subject of the training images. This highlights the value of orthogonality: while OFT also produces competitive results, LoRA and HRA struggle to generate realistic 3D icons. Moreover, HOFT and SHOFT produce accurate text in the lego sets, while the rest do not achieve it. Additional qualitative examples can be found in Appendix E.
@@ 4.4 Mathematical reasoning
For the mathematical reasoning experiments, we follow the HRA guidelines [53]. We fine-tune LLaMA2-7B [45] on the MetaMathQA dataset [52], which contains a diverse amount of mathematical questions along with rationalized answers. HOFT and SHOFT are evaluated on the GSM8K [10] and MATH [52] validation sets. Table 5 shows the accuracy of these methods alongside other PEFT baselines.
@@ 4.5 QHOFT: Quantized HOFT
In addition to the previous mathematical reasoning experiment, two additional experiments are performed in order to test the quantized versions of HOFT and SHOFT. We adapt 4-bit quantized [12] LLaMA2-7B [45] and LLaMA3.1-8B [16] to GSM8K [10] and Orca-Math [32] datasets separately and evaluate them on their respective test datasets. In particular, we follow DoRA [27] Orca-Math experimental setup: 100K elements for training and 2K for evaluation. The experimental results are reported in Table 6.
@@
The results in Table 6 demonstrate that the quantized versions of HOFT and SHOFT consistently outperform QLoRA and QDoRA under extreme parameter efficiency. On LLaMA2-7B, QHOFT achieves the highest GSM8K accuracy, while QSHOFT leads on Orca-Math. On the larger LLaMA3.1-8B model, QSHOFT delivers the best GSM8K performance, and QHOFT achieves the best Orca-Math score. These results confirm that QHOFT and QSHOFT perform well even with aggressive 4-bit quantization.
@@ A.2 Machine translation experiments
For machine translation experiments, we use a NVIDIA A30 GPU for training NLLB-3.3B model. For training LLaMA2-7B and LLaMA3.1-8B, a NVIDIA A40 GPU was used. For all experiments, the rank $r$ was set to 16, and a dropout of 0.05. The optimizer employed was AdamW with Linear LR Scheduler. For French and German datasets, models were trained for 2 epochs on the first 10K elements of the training dataset. For Slovene and Latvian, models were trained for 3 epochs. All experiments use batch size of 16 and accumulation step of 4. The number of warmup steps was set to 100. The adapted layers were Query, Key and Value. We provide in Table 8 the learning rates used per language, model and per PEFT method.
@@ A.3 Subject-driven generation experiments
For quantitative subject-driven experiments, we employ 10 NVIDIA A40 GPUs for training the Stable Diffusion 1.5 model. For all experiments, no dropout was used. The optimizer employed was AdamW with Linear LR Scheduler. All models were trained for 2005 steps using a batch size of 1. The adapted layers were Query, Key, Value and Out from the U-Net part. The learning rate used for training both HOFT and SHOFT is 5e-4.
@@ A.4 Mathematical reasoning experiments
For mathematical reasoning experiments, we employ a NVIDIA H100 GPU for training LLaMA2-7B model. For all experiments, the rank $r$ was set to 8, and no dropout. The optimizer employed was AdamW with Linear LR Scheduler. All models were trained for 2 epochs using a batch size of 8 and accumulation step of 2. The warmup ratio was set to 0.05. The adapted layers were Query and Value. The learning rates used by HOFT and SHOFT were 1e-3 and 7e-4 respectively.
@@ A.5 Mathematical reasoning experiments with quantized models
For experiments in mathematical reasoning with quantized models, we employ a NVIDIA H100 GPU for training LLaMA2-7B and LLaMA3.1-8B models. Models are quantized using NF4 and double quatization. For all experiments, the rank $r$ was set to 16, and a dropout of 0.05. The optimizer employed was AdamW with Linear LR Scheduler. For Orca-Math dataset, all models were trained for 2 epochs using a batch size of 4 and accumulation step of 1. For GSM8K dataset, all models were trained for 3 epochs using a batch size of 4 and accumulation step of 1. The number of warmup steps was set to 100. The adapted layers were Query, Key and Value. We provide in Table 9 the learning rates used per model and per PEFT method.
@@ E.1 Rank exploration
E.2 More qualitative results on subject-driven generation

### v2

@@ 4.2 Machine Translation
For measuring machine translation performance, HOFT and SHOFT are compared with DoRA, LoRA, OFT and BOFT using four languages from the CoVoST 2 (Wang et al., 2020) dataset: Slovene, German, Latvian and French. We chose these languages in order to have two well-represented languages and two low-resource languages. For French and German, models are trained on the first 10K elements of the training split. Three models are adapted for this task: NLLB-3.3B (Costa-Jussà et al., 2022), LLaMA2-7B (Touvron et al., 2023), and LLaMA3.1-8B (Grattafiori et al., 2024). We set $r=16$ for all PEFT methods and train the models for 2 epochs. Both BLEU (Papineni et al., 2002; Post, 2018) and COMET (Rei et al., 2020, 2022) results are provided for each individual language per model and PEFT method. Results obtained are shown in Table 3. We additionally provide baseline performance of the models.
From Table 3 we can observe how HOFT and SHOFT provide competitive results in French and German. In Latvian, HOFT and SHOFT give similar results in the case of NLLB-3.3B. For Slovene, both methods clearly outperform LoRA and DoRA with LLaMA2-7B, while HOFT in BLEU and SHOFT in COMET with NLLB-3.3B. Notably, the difference on both metrics is significantly higher with LLaMA3.1-8B. Compared to other orthogonal PEFT methods, such as OFT and BOFT, HOFT/SHOFT produce superior results. Overall, the top BLEU and COMET scores are almost always achieved by either HOFT or SHOFT, underlining their consistent effectiveness across multiple languages.
@@ 4.3 Subject-driven generation
For subject-driven generation, we follow the experimental protocol of HRA (Yuan et al., 2024), using the DreamBooth dataset (Ruiz et al., 2023) to train and evaluate on 25 distinct subjects, each with 30 associated prompts. We adapt the pre-trained Stable Diffusion (SD) model (Rombach et al., 2022) and compare PEFT methods quantitatively across four metrics: subject fidelity (DINO (Caron et al., 2021) and CLIP-I (Radford et al., 2021)), prompt fidelity (CLIP-T (Radford et al., 2021)), and sample diversity (LPIPS (Zhang et al., 2018)).
@@
The results, together with the provided baselines, are summarized in Table 4. Both HOFT and SHOFT outperform all baselines in subject fidelity. In terms of textual prompt fidelity, they achieve results comparable with LoRA, OFT, BOFT, and COFT. For sample diversity, they also deliver competitive performance, with the exception of BOFT in sample diversity. Additionally, we also tested HOFT and SHOFT at half the rank. Even with fewer trainable parameters, both methods consistently outperform LoRA, OFT, and COFT across all metrics, while remaining competitive with HRA on subject fidelity.
@@
Therefore, in order to gain a deeper insight into subject fidelity, we conducted an additional experiment following DoRA (Liu et al., 2024a). We fine-tuned a pre-trained Stable Diffusion XL (SDXL) model (Podell et al., 2023) on two datasets: 3D icons and lego sets. In Figure 3 we can see some examples of the styles to be learned. In this experiment, five PEFT methods are used for evaluation: LoRA, HRA, OFT, BOFT, HOFT, and SHOFT. To ensure a fair comparison, all methods used the same random sample seed for generating the images.
@@
As shown in Figure 4, HOFT and SHOFT provide better personalization than LoRA, HRA, OFT and BOFT. When generating 3D icons, both methods closely match the style and subject of the training images. This highlights the value of orthogonality: while OFT also produces competitive results, LoRA, HRA and BOFT struggle to generate realistic 3D icons. Moreover, HOFT and SHOFT produce accurate text in the lego sets, while the rest do not achieve it. Additional qualitative examples can be found in Appendix E.
@@ 4.4 Mathematical reasoning
For the mathematical reasoning experiments, we follow the HRA guidelines (Yuan et al., 2024). We fine-tune LLaMA2-7B (Touvron et al., 2023) on the MetaMathQA dataset (Yu et al., 2023), which contains a diverse amount of mathematical questions along with rationalized answers. HOFT and SHOFT are evaluated on the GSM8K (Cobbe et al., 2021) and MATH (Yu et al., 2023) validation sets. Table 5 shows the accuracy of these methods alongside other PEFT baselines.
@@ 4.5 QHOFT: Quantized HOFT
In addition to the previous mathematical reasoning experiment, two additional experiments are performed in order to test the quantized versions of HOFT and SHOFT. We adapt 4-bit quantized (Dettmers et al., 2024) LLaMA2-7B (Touvron et al., 2023), LLaMA3.1-8B (Grattafiori et al., 2024), Qwen2.5-32B (Yang et al., 2024) and LLaMA3.1-70B (Grattafiori et al., 2024) to GSM8K (Cobbe et al., 2021) and Orca-Math (Mitra et al., 2024) datasets separately and evaluate them on their respective test datasets. In particular, we follow DoRA (Liu et al., 2024a) Orca-Math experimental setup: 100K elements for training and 2K for evaluation. The experimental results are reported in Table 6.
@@
Qwen2.5-32B QLoRA 0.07 68.0 77.4
QDoRA 0.07 68.4 61.4
QHOFT 0.07 71.8 80.6
QSHOFT 0.07 75.0 66.9
LLaMA3.1-70B QLoRA 0.12 76.4 82.7
QDoRA 0.12 77.7 OOM
QHOFT 0.12 78.7 80.2
QSHOFT 0.12 70.6 81.8
The results in Table 6 demonstrate that the quantized versions of HOFT and SHOFT consistently outperform QLoRA and QDoRA under extreme parameter efficiency. On LLaMA2-7B, QHOFT achieves the highest GSM8K accuracy, while QSHOFT leads on Orca-Math. On the larger LLaMA3.1-8B model, QSHOFT delivers the best GSM8K performance, and QHOFT achieves the best Orca-Math score. These results confirm that QHOFT and QSHOFT perform well even with aggressive 4-bit quantization. Furthermore, similar trends are observed on larger models: with Qwen2.5-32B and LLaMA3.1-70B, QHOFT and QSHOFT achieve the best results on GSM8K, and QHOFT also leads on Orca-Math with Qwen2.5-32B. The only exception arises on Orca-Math with LLaMA3.1-70B, where QLoRA slightly outperforms QHOFT and QSHOFT. It is worth noting, however, that this latter experiment was conducted at the limit of our computational resources, using a batch size of 1 with 4 accumulation steps, and thus the stability of the fine-tuning process cannot be fully guaranteed.
@@ A.2 Machine translation experiments
For machine translation experiments, we use a NVIDIA A30 GPU for training NLLB-3.3B model. For training LLaMA2-7B and LLaMA3.1-8B, a NVIDIA A40 GPU was used.
@@
For all experiments, the rank $r$ was set to 16, and a dropout of 0.05. The optimizer employed was AdamW with Linear LR Scheduler. For French and German datasets, models were trained for 2 epochs on the first 10K elements of the training dataset. For Slovene and Latvian, models were trained for 3 epochs. All experiments use batch size of 16 and accumulation step of 4. The number of warmup steps was set to 100. The adapted layers were Query, Key and Value. We provide in Table 8 the learning rates used per language, model and per PEFT method.
@@ A.3 Subject-driven generation experiments
For quantitative subject-driven experiments, we employ 10 NVIDIA A40 GPUs for training the Stable Diffusion 1.5 model. For all experiments, no dropout was used. The optimizer employed was AdamW with Linear LR Scheduler. All models were trained for 2005 steps using a batch size of 1. The adapted layers were Query, Key, Value and Out from the U-Net part. The learning rate used for training BOFT, HOFT and SHOFT is 5e-4.
@@ A.4 Mathematical reasoning experiments
For mathematical reasoning experiments, we employ a NVIDIA H100 GPU for training LLaMA2-7B model. For all experiments, the rank $r$ was set to 8, and no dropout. The optimizer employed was AdamW with Linear LR Scheduler. All models were trained for 2 epochs using a batch size of 8 and accumulation step of 2. The warmup ratio was set to 0.05. The adapted layers were Query and Value. The learning rates used by DoRA, HOFT and SHOFT were 1e-3, 1e-3 and 7e-4 respectively.
@@ A.5 Mathematical reasoning experiments with quantized models
For experiments in mathematical reasoning with quantized models, we employ a NVIDIA H100 GPU for training all models.
@@
Qwen2.5-32B QLoRA 3e-4 9e-5
QDoRA 3e-4 3e-4
QHOFT 3e-4 9e-5
QSHOFT 4e-4 5e-4
LLaMA3.1-70B QLoRA 5e-4 9e-5
QDoRA 4e-4 -
QHOFT 9e-5 9e-5
QSHOFT 4e-4 9e-5
Models are quantized using NF4 and double quatization. For all experiments, the rank $r$ was set to 16, and a dropout of 0.05. The optimizer employed was AdamW with Linear LR Scheduler. For Orca-Math dataset, all models were trained for 2 epochs using a batch size of 4 and accumulation step of 1. For GSM8K dataset, all models were trained for 3 epochs using a batch size of 4 and accumulation step of 1. The number of warmup steps was set to 100. The adapted layers were Query, Key and Value. We provide in Table 9 the learning rates used per model and per PEFT method.
@@ Appendix D Time and memory consumption
Table 11: Memory and time complexity comparison on the subject-driven generation experiments
@@
Qwen2.5-32B QLoRA 8.2 52.3
QDoRA 9.8 59.7
QHOFT 9.0 53.7
QSHOFT 9.0 54.1
LLaMA3.1-70B QLoRA 7.4 62.9
QDoRA 9.6 -
QHOFT 8.1 64.3
QSHOFT 8.1 65.7
@@ E.1 Rank exploration
E.2 Training set size
To assess the robustness of HOFT and SHOFT under an increasing amount of training data, we have performed an ablation study on LLaMA3.1-8B using 5%, 10%, 25%, 50% and 100% of the commonsense reasoning training split. The results are summarized below:
Table 13: Ablation study on training data size
Method 5% 10% 25% 50% 100%
LoRA 81.8 83.5 85.0 86.4 87.0
DoRA 81.7 83.4 84.9 86.5 87.1
OFT 77.8 80.7 83.2 85.0 86.5
HOFT 82.0 83.7 85.1 86.5 87.4
SHOFT 82.3 83.6 85.4 86.3 87.1
As shown, both HOFT and SHOFT consistently surpass other PEFT methods across all data regimes, with SHOFT leading at the lowest-data setting (5%) and HOFT achieving the highest accuracy at the full-data scale (100%). We will include the complete experimental protocol, hyperparameter settings, and detailed results in the revised manuscript.
E.3 More qualitative results on subject-driven generation